Assume a member is selected at random from the population represented by the graph. Find the probability that the member seleA
Answer #1

ANSWER Griven M = 507 E116 XI-M 2,5 200 -507 1) 116 2.0465 22= 475-507 TIG -0.2758 PC 200<x£475)= P(-2.6465<2< -0.2758) 1) p